OpenStudy (anonymous):

What is the description of y=sqrt(9x-3) to make it easy to graph using transformations of its parent function?

OpenStudy (anonymous):

A) the graph of y=3sqrt(x), shifted right 3 units B)the graph of y=3sqrt(x), shifted right 1/3 unit C) The graph of y=sqrt(x), shifted right 3 units and up 9 units D) The graph of y=sqrt(x), shiftes right 1/3 unit and up 9 units Those are my choices :)

OpenStudy (anonymous):

@callisto

OpenStudy (anonymous):

@actionsports do you know anything on this?

OpenStudy (anonymous):

the answer is B - the graph of y=3 square root of x, shifted right 1/3 unit

OpenStudy (anonymous):

Thank you so much!

OpenStudy (k_lynn):

i have no idea how to do this yet...maybe @sammixboo @iGreen or @confluxepic can help you.

OpenStudy (anonymous):

no prob:)

OpenStudy (anonymous):

Can you help me with another quick one @actionsports ?

OpenStudy (anonymous):

sure:)

OpenStudy (confluxepic):

@k_lynn The question is already solved.

OpenStudy (k_lynn):

oops, sorry.